OpenStudy (anonymous):

A motor car of mass 1000kg travelling at 90 km per hour is brought to rest by the brakes in 100m. Calculate the initial kinetic energy and the average breaking force required. URGENT PLEASE

OpenStudy (anonymous):

m=1000 kg, V= 90 km/hr= 25m/s brought to rest in 100m now \[U^2=2*a*100\] and \[(25)^2=2*a*100----a=\frac{ 625 }{ 200 }=6.25m/s^2\] so average force = m*a=1000*6.25=6250N

OpenStudy (anonymous):

and initial K.E= 0.5 m*v^2 that is \[K.E=\frac{ 1 }{ 2 }*m*V^2=0.5*1000*(25)^2=312500 J= 3.125*10^5 J\]